Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
Technical integration and coordination SASE1 Status and SASE3 Readiness
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 594     Entry time: 10 Mar 2025, 09:53
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 10.03.2025 
General Announcement:
- From the DRC side, there is nothing particular to report.
- From the PRC side, nothing apart of what was reported below.
- XFELO studies: J. Gruenert says a successful time last week to bring all diagnostics in operation; for the first time full system was operational and detected; length of the cavity was measured; 
it was a major step forward by having a signal that can be tuned or amplified.

CTRL:
- Nothing to report.

DA:
- Not present.

EEE:
- Nothing to report.

VAC:
- Friday issue with ion pump controller in XTD1: exchanged by VAC OCD and EEE colleagues in a ZZ.
- Sunday worked in XTD6 on the PBLM - tunnel is open and will be closed after lunch.
- Work ongoing to check piezo at the monochromators.
- No news on the Mono vacuum issue last week at HED (reason).

XPD:
- Blue week studies were successful (e.g. XFELO diagnostics (adjusted grating by ZZ on Friday))

DET:
- No major issue to report.
- SPB pre-intervention with AGIPD took place last week: replaced module and tested - at the moment module is operational; 
however, connection was lost and needs time to evaluate if it will work in stable conditions. Intervention in summer will be later decided.

ITDM: 
- Nothing to report.

XRO:
- During tunnel activity on Sunday also a temperature sensor was exchanged at the PBLM.

LAS:
- Not present.

UNSYS:
- Undulators are all operational.
- Apple-X studies last week: very successful; it reached rotation of microbunching, beam separation and measured contrast (90); 
it was found the orbit before SA3 was not optimal and explain why the expected contrast was not reached; due to ZZ could not use full beamtime allocation; needs further development to improve contrast. 

SPB:
- Pulse arrival time tests were done; an issue needs to be corrected, in touch with experts. 

FXE:
- Issue with vacuum interlock (3 ions pumps): was a PLC issue resolved by EEE.

MID:
- Nothing to report.

HED:
- Nothing to report.

SCS:
- Nothing to report.

SQS:
- Nothing to report.

SXP:
- Commissioning last week went well; GATT could not be set on Friday; R. Villanueva comments the interlock is expected to be set and is a feature to avoid sudden jumps in pressure. 
After stopping and restarting injection, this releases the interlock.
ELOG V3.1.4-7c3fd00